# Learn by Reasoning: Analogical Weight Generation for Few-Shot Class-Incremental Learning

### v1

@@ III-C2 Weight Self Attention (WSA)
$Q_{s}={\boldsymbol{q}_{W}}+\boldsymbol{d}_{E},K_{s}={\boldsymbol{q}_{W}}+\boldsymbol{d}_{E},V_{s}=\boldsymbol{d}_{E}$ (8)
@@
where $D$ is the dimension of the decoder embedding. WSA enhances the efficiency of knowledge utilization by performing self-attention processing on $\boldsymbol{d}_{E}$ and $\boldsymbol{q}_{W}$, supplementing the knowledge related to new class weights. The supplementary knowledge generated by WSA aids in efficiently learning new classes with minimal data, providing a more comprehensive understanding of the new classes.
TABLE I: Performance of FSCIL in each session on miniImageNet and comparison with other studies. The results were implemented by [1, 5] in the FSCIL setting. “Average Acc.” is the average accuracy of all sessions. “Average Improv.” indicates the overall improvement of BiAG over each method’s average accuracy. “Final Improv.” calculates the improvement of our method in the last session.
@@
CLOM (NeurIPS-2022) 73.08 68.09 64.16 60.41 57.41 54.29 51.54 49.37 48.00 58.48 +11.57 +11.57
MetaFSCI (CVPR-2022) 72.04 67.94 63.77 60.29 57.58 55.16 52.90 50.79 49.19 58.85 +11.20 +11.20
LIMIT (ECCV-2022) 72.32 68.47 64.30 60.78 57.95 55.07 52.70 50.72 49.19 59.06 +10.99 +10.99
FACT (CVPR-2022) 74.60 72.09 67.56 63.52 61.38 58.36 56.28 54.24 52.10 62.24 +7.81 +7.73
@@
III-C3 Weight & Proto Analogical Attention (WPAA)
@@
${\boldsymbol{{q}_{P}}={\varphi}_{\rm{{}_{SMC}}}(\boldsymbol{{q}_{L}})},$ (10)
@@
$\rm{W}^{t}=\rm{softmax}\Big{(}\frac{{{Q}_{c}}\cdot K_{c}^{T}}{\sqrt{D}}\Big{)}\cdot V_{c}.$ (15)
@@
For scalability, WPAA can be enhanced with SCM, which converts the previous layer’s output into class prototype semantics and adds them to the learnable query for updates:
$\boldsymbol{{q}_{L_{n}}}={\varphi}_{\rm{{}_{SCM}}}(\rm{W}^{t})+\boldsymbol{{q}_{L_{(}n-1)}}$ (16)
Where n is the number of layers in BiAG. The WSA and WPAA parameters in each layer are independent, while SCM shares parameters across layers. This design ensures consistent semantic conversion and enhances the model’s analogy capability through layer-by-layer optimization.
IV Experiments
IV-A Datasets
CIFAR-100 consists of 100 general object categories with relatively low image resolution. In our setting, the base training set $\mathcal{D}^{0}$ contains 60 labeled classes, which are used for initial supervised training. The remaining 40 classes are divided into incremental sessions. In each session $t>0$, 5 new classes are introduced with only 5 labeled samples per class, following a 5-way 5-shot format.
miniImageNet also includes 100 image categories and serves as a popular benchmark for few-shot learning. Similar to CIFAR-100, we adopt 60 classes as the base set $\mathcal{D}^{0}$ and use the remaining 40 classes for incremental learning. Each incremental session adds 5 new classes, with 5 labeled samples per class, maintaining the 5-way 5-shot configuration.
TABLE II: Performance of FSCIL in each session on CUB200 and comparison with other studies.
@@ IV-A Datasets
KRRM (TCSVT-2024) 79.46 76.11 73.12 69.31 67.97 65.86 64.50 63.83 62.20 62.00 60.97 67.85 +2.86 +2.75
@@
SPPR (NeurIPS-2021) 64.10 65.86 61.36 57.45 53.69 50.75 48.58 45.66 43.25 54.52 +14.41 +14.62
LIMIT (ECCV-2022) 74.30 70.44 66.12 61.24 58.77 53.91 51.36 49.20 45.33 58.63 +10.30 +12.62
C-FSCIL (CVPR-2022) 77.47 72.00 66.54 61.39 58.10 55.12 52.47 50.47 47.11 60.74 +8.19 +10.84
Data-free Replay (CVPR-2022) 74.40 70.60 66.54 61.39 58.91 51.36 50.14 47.76 44.10 58.18 +10.75 +13.85
MetaFSCIL (CVPR-2022) 74.50 70.10 66.87 62.77 59.48 56.52 54.32 52.56 49.97 60.79 +8.14 +8.98
@@
ALFSCIL (TCSVT-2024) 80.75 77.88 72.94 68.79 65.33 62.15 60.02 57.68 55.17 65.81 +3.12 +2.78
KRRM (TCSVT-2024) 81.25 77.23 73.30 69.41 66.69 63.93 62.16 59.62 57.41 66.00 +2.93 +0.54
@@
ALICE (ECCV-2022) 79.00 56.63 64.30
FACT (CVPR-2022) 75.03 53.71 62.97
@@
BiDist (CVPR-2023) 78.41 55.82 64.80
ALFSCIL (TCSVT-2024) 80.60 55.93 65.91
SVAC (CVPR-2023) 80.58 57.57 66.68
KRRM (TCSVT-2024) 81.12 58.91 67.29
@@ IV-B Implementation Details
For training in the base session, we run 200 epochs on both CIFAR-100 and miniImageNet, using an initial learning rate of 0.1 and a batch size of 128. For the CUB-200 dataset, which is more fine-grained and prone to overfitting, we set the initial learning rate to 0.02 and reduced the batch size to 64 to stabilize training. The learning rate is decayed by a factor of 0.1 at 100 and 150 epochs. All models are trained on a single NVIDIA 3090 GPU using PyTorch. This consistent and well-established training setup ensures a fair comparison with prior FSCIL methods while allowing us to focus on evaluating the impact of our proposed analogical learning framework.
@@ IV-C Comparison with State-of-the-Art Methods
On miniImageNet, as presented in Table I, our method achieves the highest final session accuracy of 59.83% and the best average accuracy of 70.05%. Compared to Yourself and NC-FSCIL, two of the strongest recent methods, BiAG achieves improvements of 1.25% and 2.01% in final accuracy and 0.83% and 2.23% in average accuracy, respectively. These consistent gains demonstrate the effectiveness of our analogical reasoning framework in maintaining discriminative power across incremental sessions, especially in the later stages where performance often degrades.
On CUB-200, a fine-grained dataset with high intra-class similarity, our method achieves a final session accuracy of 63.72% and an average accuracy of 70.71%, as detailed in Table II. Compared to SVAC, BiAG shows an improvement of 1.36% in final accuracy and 1.23% in average accuracy. Notably, BiAG also surpasses Yourself, indicating its capacity to handle fine-grained incremental learning with stronger semantic consistency. These results highlight that our method not only performs well on general datasets but also excels under more challenging fine-grained scenarios.
@@
Across all three benchmarks, we report the average base session accuracy, final session accuracy, and overall average accuracy to comprehensively evaluate the performance of different methods under the FSCIL setting. As summarized in Table IV, our method BiAG consistently achieves the best results across all three metrics. In terms of Average final session accuracy, which reflects the model’s ability to retain base knowledge while integrating new classes over time, BiAG achieves 60.50%. This clearly exceeds Yourself and KRRM, suggesting that our method suffers less from catastrophic forgetting and is more stable throughout incremental sessions. For the overall average accuracy, BiAG again achieves the highest score of 69.90%, demonstrating superior performance across all sessions. This result surpasses both Yourself and NC-FSCIL, achieving relative improvements of 1.36% and 2.37%, respectively. These consistent improvements across all metrics highlight the robustness, generalization, and effectiveness of the proposed analogical reasoning framework. BiAG is not only capable of building strong initial representations but also excels in retaining knowledge and adapting to new categories in both general and fine-grained FSCIL scenarios.
TABLE V: Performance of base and new classes on three datasets.
Method Final Session Final Session Final Session
Accuracy Base Acc. New Avg Acc.
CEC 49.35 68.78 25.17
SVAC 57.57 75.12 33.79
NC-FSCIL 57.95 75.49 34.44
BiAG (Ours) 60.50 78.34 36.21
TABLE VI: Comparison of final session base and new class accuracy on CIFAR-100. ”-” indicates results not reported in [41, 42].
@@
TABLE VII: Comparison of final session base and new class accuracy on miniImageNet. ”-” indicates results not reported in [41, 42].
@@ IV-D Performance on Base and Novel Classes
TABLE VIII: Comparison of Last Session base classes accuracy, Last Session new classes accuracy, Last Session new classes average accuracy, and average accuracy on CUB-200. ”-”: Results not reported in OrCo [51]
Acc. Base Acc. New Avg Acc. New Acc. Acc.
CEC (CVPR-2021) 52.12 70.46 33.78 34.23 61.33
FACT (CVPR-2022) 56.94 73.90 39.98 40.50 62.42
LIMIT (ECCV-2022) 57.41 73.60 41.22 41.80 66.67
ALFSCIL (TCSVT-2024) 59.30 74.21 44.39 45.17 66.99
NC-FSCIL (ICLR-2023) 59.44 73.98 44.90 45.83 67.28
OrCo (CVPR-2024) - 66.62 - 49.25 62.36
SVAC (CVPR-2023) 62.50 77.65 47.35 47.68 69.61
BiAG (Ours) 63.72 78.57 48.87 49.36 70.71
@@
As shown in Table V, BiAG achieves the best results across all three datasets, demonstrating superior overall performance and balance. It reaches 60.50% final session accuracy, 78.34% base accuracy, and 36.21% new class average accuracy—outperforming NC-FSCIL by 2.55% on base accuracy and 1.77% on new average accuracy. These results highlight that BiAG provides a robust and scalable framework for FSCIL, exhibiting consistently strong performance in both knowledge retention and novel class generalization. Its effectiveness across diverse datasets and evaluation metrics validates the strength of our analogical learning strategy.
@@
On CIFAR-100 (Table VI), BiAG achieves the best performance across all three metrics—final session accuracy, base class accuracy, and novel class average accuracy. It outperforms NC-FSCIL in every aspect, demonstrating a superior balance between knowledge retention and adaptation and validating its robustness. On miniImageNet (Table VII), BiAG also achieves the highest overall and base class accuracy, surpassing NC-FSCIL by 3.28% in base retention.
@@
On the fine-grained benchmark CUB-200 (Table VIII), BiAG achieves SOTA performance across all evaluation metrics. Compared with SVAC, one of the strongest prior methods, BiAG yields a clear improvement of 1.22% in final session base class accuracy and 1.52% in the average accuracy across all novel classes in the final session. In the final 5-way incremental step, BiAG achieves 49.36% accuracy, exceeding OrCo, which focuses more on improving performance for the currently added novel classes. These results underline BiAG’s effectiveness in fine-grained FSCIL settings, where classes often exhibit high visual similarity. The consistent improvements across both old and new classes suggest that BiAG not only retains previously learned knowledge but also excels at distinguishing subtle differences among closely related new categories.
Figure 4: The results of ablation experiments.Ablation studies on the proposed BiAG. (a) Final session accuracy on three datasets after incrementally adding key components (WPAA, WSA, SMC); (b) Session-wise accuracy comparison on CUB-200, showing the contribution of each component throughout the incremental process; (c) Accuracy comparison of BiAG with different numbers of layers, demonstrating that a deeper design improves generalization to new classes.
TABLE IX: Ablation study on CIFAR-100 showing the individual and combined effects of the WPAA, WSA, and SMC modules within BiAG. We report the accuracy for each session, the overall average accuracy (Average ACC), the improvement over the baseline (Average Improv), and the accuracy gain in the final session (Final Improv.).
WPAA WSA SMC Acc. in each session (%) $\uparrow$ Average Average Final
0 1 2 3 4 5 6 7 8 Acc. Improv. Improv.
84.00 65.32 63.24 61.84 59.59 58.24 56.80 55.97 54.92 62.21
84.00 70.32 67.47 65.25 62.94 61.82 60.60 58.98 57.60 65.44 +3.23 +2.68
84.00 71.83 70.24 66.75 64.61 63.48 61.24 60.41 57.73 66.70 +4.49 +2.81
84.00 78.97 74.73 70.75 67.36 64.21 62.21 60.20 57.95 68.93 +6.72 +3.03
IV-E Ablation Study

### v2

@@ III-C2 Weight Self-Attention (WSA)
$Q_{s}=K_{s}=\begin{cases}q_{W}+d_{E},&n=1,\\ q_{W}+\rm{W}^{\textit{t}}_{\textit{n-1}},&n>1,\end{cases}$ (9)
$V_{s}=\begin{cases}d_{E},&n=1,\\ \rm{W}^{\textit{t}}_{\textit{n-1}},&n>1,\end{cases}$ (10)
@@
where $D$ is the dimension of the decoder embedding and $n$ denotes the layer index in the stacked BiAG block (with n=1n{=}1 being the first layer). WSA enhances the efficiency of knowledge utilization by performing self-attention on $\boldsymbol{d}_{E}$ and $\boldsymbol{q}_{W}$, thereby supplementing knowledge related to new class weights. Here, $d_{E}$ denotes the decoder embedding in BiAG. It is trained only during the Analogical Generator Training stage to encode old-class knowledge as a fixed, session-invariant prior and to enable forward-only incremental updates; in all incremental sessions $d_{E}$ is frozen. At the first layer, $d_{E}$ is initialized to $\mathbf{0}$ before training to avoid interference from other information; for deeper layers, $d_{E}$ is taken as the output of the previous layer.
In this design, both $Q_{s}$ and $K_{s}$ incorporate the SCM-transformed query $q_{W}$, ensuring that attention is explicitly guided by new-class semantics, while the value path $V_{s}$ remains free of $q_{L}$ to preserve a stable memory content across layers. This separation allows each layer to refine the representation progressively without contaminating the inter-layer content flow.
WSA therefore serves as a preparatory stage in our analogical reasoning framework: it refines the new-class representation by emphasizing salient attributes and aligning them with prior knowledge, akin to how humans first identify the key features of a new concept before drawing analogies to known categories. By augmenting prototypes with supplementary weight-level knowledge, WSA enables new classes to be efficiently learned from minimal data and yields a more complete and semantically expressive representation, thereby providing a stronger foundation for subsequent analogical reasoning in WPAA.
TABLE I: Performance of FSCIL in each session on miniImageNet. The results were implemented by [1, 4] in the FSCIL setting. “Average Acc.” is the average accuracy of all sessions. “Average Improv.” indicates the overall improvement of BiAG over each method’s average accuracy. “Final Improv.” calculates the improvement of our method in the last session.
@@
CLOM (NeurIPS-2022) 73.08 68.09 64.16 60.41 57.41 54.29 51.54 49.37 48.00 58.48 +11.57 +11.83
MetaFSCIL (CVPR-2022) 72.04 67.94 63.77 60.29 57.58 55.16 52.90 50.79 49.19 58.85 +11.20 +10.64
LIMIT (ECCV-2022) 72.32 68.47 64.30 60.78 57.95 55.07 52.70 50.72 49.19 59.06 +10.99 +10.64
FACT (CVPR-2022) 72.56 69.63 66.38 62.77 60.60 57.33 54.34 52.16 50.49 60.69 +9.36 +9.34
@@
III-C3 Weight & Prototype Analogical Attention (WPAA)
@@
${\boldsymbol{{q}_{P}}={\varphi}_{\rm{{}_{{SCM}}}}(\boldsymbol{{q}_{L}})},$ (12)
@@
$\rm{W}^{t}=\rm{softmax}\Big(\frac{{{Q}_{c}}\cdot K_{c}^{T}}{\sqrt{D}}\Big)\cdot{\textit{V}}_{c}.$ (17)
@@
For scalability, we design WPAA to work together with the SCM. Specifically, SCM converts the WPAA output from the previous layer from weight semantics into prototype semantics, and then adds it to the learnable query from the (n1)(n\!-\!1)-th layer to form the query for the current $n$-th layer:
$\boldsymbol{{q}_{L_{n}}}={\varphi}_{\rm{{}_{SCM}}}(\rm{W}^{t}_{n-1})+\boldsymbol{{q}_{L_{(}n-1)}}\$(18)
Here, n denotes the index of the current layer in BiAG, and $W^{t}_{\,n-1}$ is the WPAA output produced in the previous layer
The parameters of WSA and WPAA are independent across different layers, while SCM shares its parameters globally. This design ensures semantic consistency across layers and progressively refines the learnable query, thus enhancing the model’s analogical reasoning ability through iterative layer-by-layer optimization.
TABLE II: Performance of FSCIL in each session on CUB-200 and comparison with other studies.
@@
KRRM (TCSVT-2024) 79.46 76.11 73.12 69.31 67.97 65.86 64.50 63.83 62.20 62.00 60.97 67.76 +2.95 +2.75
Comp-FSCIL (ICML-2024) 80.94 77.51 74.34 71.00 68.77 66.41 64.85 63.92 62.12 62.10 61.17 68.47 +2.24 +2.55
@@
IV Experiments
IV-A Datasets
CIFAR-100 consists of 100 general object classes with relatively low image resolution. In our setting, the base training set $\mathcal{D}^{0}$ contains 60 labeled classes, which are used for initial supervised training. The remaining 40 classes are divided into incremental sessions. In each session $t>0$, 5 new classes are introduced with only 5 labeled samples per class, following a 5-way 5-shot format.
miniImageNet also includes 100 image classes and serves as a popular benchmark for few-shot learning. Similar to CIFAR-100, we adopt 60 classes as the base set $\mathcal{D}^{0}$ and use the remaining 40 classes for incremental learning. Each incremental session adds 5 new classes, with 5 labeled samples per class, maintaining the 5-way 5-shot configuration.
@@ IV-A Datasets
SPPR (NeurIPS-2021) 64.10 65.86 61.36 57.45 53.69 50.75 48.58 45.66 43.25 54.52 +14.41 +14.70
LIMIT (ECCV-2022) 74.30 70.44 66.12 61.24 58.77 53.91 51.36 49.20 45.33 58.96 +9.97 +12.62
C-FSCIL (CVPR-2022) 77.47 72.00 66.54 61.39 58.10 55.12 52.47 50.47 47.11 60.07 +8.86 +10.84
Data-free Replay (ECCV-2022) 74.40 70.20 66.54 62.51 59.71 56.58 54.52 52.39 50.14 60.78 +8.15 +7.81
MetaFSCIL (CVPR-2022) 74.50 70.10 66.87 62.77 59.48 56.52 54.32 52.56 49.97 60.79 +8.14 +7.98
@@
ALFSCIL (TCSVT-2024) 80.75 77.88 72.94 68.79 65.33 62.15 60.02 57.68 55.17 66.75 +2.18 +2.78
KRRM (TCSVT-2024) 81.25 77.23 73.30 69.41 66.69 63.93 62.16 59.62 57.41 67.89 +1.04 +0.54
Comp-FSCIL(ICML-2024) 80.93 76.52 72.69 68.52 65.50 62.62 60.96 59.27 56.71 67.08 +1.85 +1.24
@@
ALICE (ECCV-2022) 79.00 56.63 64.32
FACT (CVPR-2022) 74.35 53.18 62.45
@@
KRRM (TCSVT-2024) 81.12 58.91 68.10
@@ IV-B Implementation Details
For training in the base session, we run 200 epochs on both CIFAR-100 and miniImageNet, using an initial learning rate of 0.1 and a batch size of 128. For the CUB-200 dataset, which is more fine-grained and prone to overfitting, we set the initial learning rate to 0.02 and reduced the batch size to 64 to stabilize training. The learning rate is decayed by a factor of 0.1 at 100 and 150 epochs. All models are trained on a single NVIDIA 3090 GPU using PyTorch.
Following the base session, the analogical generator (BiAG) training stage is conducted for another 200 epochs, where the backbone is frozen and only the BiAG modules are optimized. This stage has substantially lower computational cost per epoch due to the fixed feature extractor, yet ensures sufficient adaptation of the semantic-to-weight mapping before entering the incremental sessions.
This consistent and well-established training setup ensures a fair comparison with prior FSCIL methods while allowing us to focus on evaluating the impact of our proposed analogical learning framework.
@@ IV-C Comparison with State-of-the-Art Methods
On miniImageNet, as presented in Table I, our method achieves the highest final session accuracy of 59.83% and the best average accuracy of 70.05%. Compared to YourSelf and NC-FSCIL, two of the strongest recent methods, BiAG achieves improvements of 0.83% and 1.52% in final accuracy and 1.25% and 2.23% in average accuracy, respectively. These consistent gains demonstrate the effectiveness of our analogical reasoning framework in maintaining discriminative power across incremental sessions, especially in the later stages where performance often degrades.
On CUB-200, a fine-grained dataset with high intra-class similarity, our method achieves a final session accuracy of 63.72% and an average accuracy of 70.71%, as detailed in Table II. Compared to SVAC, BiAG shows an improvement of 1.23% in final accuracy and 1.36% in average accuracy. Notably, BiAG also surpasses YourSelf, indicating its capacity to handle fine-grained incremental learning with stronger semantic consistency. These results highlight that our method not only performs well on general datasets but also excels under more challenging fine-grained scenarios.
@@
On ImageNet-1K. We evaluate BiAG on the large-scale ImageNet-1K dataset using a 50-way, 10-shot incremental setting with 500 randomly sampled base classes. As shown in Table V, BiAG surpasses SoftNet and SVAC by 6.04% and 1.61% in Average Accuracy, demonstrating strong scalability and analogical reasoning capability in large-scale FSCIL.
TABLE V: Performance on ImageNet-1K.
Method Base Acc. Final ACC. Avg Acc.
SoftNet (ICLR-2023) 36.21 21.21 27.51
SVAC (CVPR-2023) 37.74 26.12 31.94
BiAG (Ours) 37.98 28.09 33.55
TABLE VI: Comparison of Last Session base classes accuracy, Last Session new classes accuracy, Last Session new classes average accuracy, and average accuracy on CUB-200. ”-”: Results not reported in OrCo [49]
Acc. Base Acc. New Avg Acc. New Acc. Acc.
CEC (CVPR-2021) 52.12 70.46 33.78 34.23 61.33
FACT (CVPR-2022) 56.94 73.90 39.98 40.50 62.42
LIMIT (ECCV-2022) 57.41 73.60 41.22 41.80 66.67
ALFSCIL (TCSVT-2024) 59.30 74.21 44.39 45.17 66.99
NC-FSCIL (ICLR-2023) 59.44 76.19 42.69 45.83 67.28
OrCo (CVPR-2024) - 66.62 - 49.25 62.36
SVAC (CVPR-2023) 62.49 77.65 47.35 47.68 69.35
BiAG (Ours) 63.72 78.57 48.87 49.36 70.71
TABLE VII: Comparison of final session base and new class accuracy on CIFAR-100. ”-” indicates results not reported in [6, 41].
@@
TABLE VIII: Comparison of final session base and new class accuracy on miniImageNet. ”-” indicates results not reported in [6, 41].
@@
Across three benchmarks (miniImageNet, CUB-200 and CIFAR-100), we report the average base session accuracy, final session accuracy, and overall average accuracy to comprehensively evaluate the performance of different methods under the FSCIL setting. As summarized in Table IV, our method BiAG consistently achieves the best results across all three metrics. In terms of Average final session accuracy, which reflects the model’s ability to retain base knowledge while integrating new classes over time, BiAG achieves 60.50%. This clearly exceeds YourSelf and Comp-FSCIL, suggesting that our method suffers less from catastrophic forgetting and is more stable throughout incremental sessions. For the overall average accuracy, BiAG again achieves the highest score of 69.90%, demonstrating superior performance across all sessions. This result surpasses both YourSelf and Comp-FSCIL, achieving relative improvements of 1.36% and 1.77%, respectively. These consistent improvements across all metrics highlight the robustness, generalization, and effectiveness of the proposed analogical reasoning framework. BiAG is not only capable of building strong initial representations but also excels in retaining knowledge and adapting to new classes in both general and fine-grained FSCIL scenarios.
@@ IV-D Performance on Base and Novel Classes
TABLE IX: Performance of base and new classes on three datasets.
Method Final Session Final Session Final Session
Accuracy Base Acc. New Avg Acc.
CEC 49.63 68.78 27.35
SVAC 57.57 75.12 33.79
NC-FSCIL 57.95 75.49 34.44
BiAG (Ours) 60.50 78.34 36.21
TABLE X: Ablation study on CIFAR-100 showing the individual and combined effects of the WPAA, WSA, and SCM modules within BiAG. We report the accuracy for each session, the overall average accuracy (Average ACC), the improvement over the baseline (Average Improv), and the accuracy gain in the final session (Final Improv.).
WPAA WSA SCM Acc. in each session (%) $\uparrow$ Average Average Final
0 1 2 3 4 5 6 7 8 Acc. Improv. Improv.
84.00 67.32 63.24 61.84 59.59 58.24 56.80 55.97 54.92 62.44
84.00 70.32 64.47 65.25 62.94 61.82 60.60 58.98 57.60 65.11 +2.67 +2.68
84.00 73.83 71.24 68.75 64.61 63.48 61.24 60.01 57.73 67.21 +4.77 +2.81
84.00 78.97 74.73 70.75 67.36 64.21 62.21 60.20 57.95 68.93 +6.49 +3.03
@@
As shown in Table IX, BiAG achieves the best results across all three datasets, demonstrating superior overall performance and balance. It reaches 60.50% final session accuracy, 78.34% base accuracy, and 36.21% new class average accuracy, outperforming NC-FSCIL by 2.85% on base accuracy and 1.77% on new average accuracy. These results highlight that BiAG provides a robust and scalable framework for FSCIL, exhibiting consistently strong performance in both knowledge retention and novel class generalization. Its effectiveness across diverse datasets and evaluation metrics validates the strength of our analogical learning strategy.
@@
On CIFAR-100 (Table VII), BiAG achieves the best performance across all three metrics—final session accuracy, base class accuracy, and novel class average accuracy. It outperforms NC-FSCIL in every aspect, demonstrating a superior balance between knowledge retention and adaptation and validating its robustness. On miniImageNet (Table VIII), BiAG also achieves the highest overall and base class accuracy, surpassing NC-FSCIL by 3.28% in base retention.
@@
On the CUB-200 (Table VI), BiAG achieves SOTA performance across all evaluation metrics. Compared with SVAC, one of the strongest prior methods, BiAG yields a clear improvement of 0.92% in final session base class accuracy and 1.52% in the average accuracy across all novel classes in the final session. In the final 5-way incremental step, BiAG achieves 49.36% accuracy, exceeding OrCo, which focuses more on improving performance for the currently added novel classes. These results underline BiAG’s effectiveness in fine-grained FSCIL settings, where classes often exhibit high visual similarity. The consistent improvements across both old and new classes suggest that BiAG not only retains previously learned knowledge but also excels at distinguishing subtle differences among closely related new classes.
IV-E Comparison with CLIP-based Methods
TABLE XI: Performance comparison of our method with recent CLIP-based FSCIL methods on miniImageNet and CIFAR-100. Results are average accuracies across all sessions.
Method Backbone miniImageNet CIFAR-100
CLIP (0-shot) Vis + Text 62.67 32.24
LRT (TPAMI-2025) Vis + Text 75.94 71.50
PCL (TCSVT-2025) Vis + Text 94.02 82.92
BiAG (Ours) Vis Only 86.43 84.92
Table XI compares our BiAG method with representative CLIP-based methods on miniImageNet and CIFAR-100. CLIP-based methods (CLIP 0-shot, LRT, and PCL) utilize both visual and text encoders, whereas our BiAG relies solely on the visual encoder without any language guidance. The reported values correspond to the average accuracy across all incremental sessions under the standard FSCIL evaluation protocol. Despite this, BiAG achieves 84.92% on CIFAR-100, surpassing PCL and LRT, and performs competitively on miniImageNet. This highlights that our analogical weight generation framework provides strong generalization even without leveraging large-scale language models.
IV-F Ablation Study